# Bird-Interact: Re-imagining Text-to-SQL Evaluation for Large Language Models via Lens of Dynamic Interactions

### v2

@@ Acknowledgement
We would like to express our sincere gratitude to Irina Saparina, Mohammadreza Pourreza, Mehdi Bouzouina, Hailong Li, Jiatong Shi, and Professor Shinji Watanabe for their fruitful discussions and valuable insights that helped improve this work.
@@ Appendix A Limitations
Appendix B Annotation Group Details
@@
B.1 Annotator Entrance Test
@@ B.3 Qualification Test
Appendix C Benchmark Design Principles
@@
Appendix D Comparison with Related Benchmarks
D.1 Task Comparison
@@ D.2 Database Comparison
Appendix E Evaluation Metrics
E.1 Success Rate (SR)
@@
where $\mathbb{I}[\cdot]$ is the indicator function that equals 1 if the prediction is correct and 0 otherwise. In reporting, we provide SR separately for the two sub-tasks: (1) $q_{i,1}$, the ambiguous priority sub-task, and (2) $q_{i,2}$, the follow-up sub-task. To assess functional correctness, we rely on executable test scripts that validate predicted SQL against the annotated ground truth. Details of the test scripts are provided in Appendix F.
E.2 Normalized Reward
@@
Appendix F Test Scripts
@@ F.2 DM Queries
Appendix G Ambiguity and Follow-up Annotation Details
G.1 User Query Ambiguity Annotation
@@
For each category, we provide annotators with a structured taxonomy including type definitions, annotation conditions, and examples, ensuring systematic and consistent ambiguity injection, as outlined in Appendix G.4.
@@
G.2 Knowledge and Environmental Ambiguity Annotation
@@ G.3 Ambiguity Chain
G.4 User Query Ambiguity Taxonomy
@@
Intent-level ambiguity arises directly from user language, where the request is vague, underspecified, or missing critical details (e.g., “find elderly people” without defining the age threshold). If not resolved, intent-level ambiguity can severely degrade user experience and lead to erroneous SQL (Saparina & Lapata, 2024; Li et al., 2025f). We summarize six types of user query ambiguity in Table 6, according related works (Saparina & Lapata, 2024; Li et al., 2025f; Dong et al., 2025; Wang et al., 2020; Min et al., 2020; Bhaskar et al., 2023; Huo et al., 2024; Floratou et al., 2024; Huang et al., 2023; Ding et al., 2025) and guide the annotators to inject them into unambiguous user queries: (1) Lexical Ambiguity from tokens with multiple meanings, (2) Syntactic Ambiguity from multiple valid grammatical structures, (3) Semantic Ambiguity from vague phrasing (e.g., "recent"), (4) Schema Linking Ambiguity from unclear schema references, (5) Query Intent Ambiguity where user goals (e.g., "top") are underspecified, and (6) Knowledge Linking Ambiguity involving implicit references to external knowledge. The performance of different types is shown in Figure 14.
@@
G.5 Follow-up Sub-Task Taxonomy
@@
Appendix H Experiment Details
H.1 Choice of PostgreSQL as the Evaluation Database System
@@ H.3 Experiment Setup
Appendix I Action Space and Selection Patterns in $a$-Interact
@@
I.1 Action Space in $a$-Interact
@@
I.2 Universal Cost Scheme for Custom Agents
@@
I.3 Action Selection Patterns and Their Impact (Full Set)
@@
Appendix J Performance on Different Ambiguity Types
@@
Appendix K Experiments on Bird-Interact-Lite
@@
Model Priority Question (Success Rate %) $\uparrow$ Follow Ups (Success Rate %) $\uparrow$ Reward* $\uparrow$
Avg.
Cost $\downarrow$
@@
Gemini-2.0-Flash 8.21 44.76 21.00 4.10 21.90 10.33 17.80 0.03 \$
DeepSeek-R1 6.67 47.62 21.00 3.59 28.57 12.33 18.40 0.09 \$
GPT-4o 12.31 43.81 23.33 4.62 17.14 9.00 19.03 0.46 \$
DeepSeek-V3 11.79 44.76 23.33 6.15 16.19 9.67 19.23 0.06 \$
Qwen-3 7.18 49.52 22.00 5.64 29.52 14.00 19.60 0.03 \$
O3-Mini 14.87 45.71 25.67 6.67 21.90 12.00 21.57 0.08 \$
Claude-Sonnet-3.7 22.05 56.19 34.00 10.77 30.48 17.67 29.10 0.67 \$
@@
Appendix L Error Analysis
@@
Appendix M User Simulator Design Details
@@
Appendix N Evaluating the Function-Driven User Simulator
@@
N.1 UserSim-Guard: A Benchmark for Simulator Robustness
To enable a systematic evaluation of simulator performance, we constructed UserSim-Guard, a manually curated dataset containing 1,989 challenging questions.
@@
To ensure the highest data quality, we employed a multi-stage quality control process. Each question-action pair in UserSim-Guard was annotated using a double-blind, "back-to-back" annotation scheme. Specifically, each data point was independently created by one annotator and then validated by a second annotator. Any disagreements between the two annotators were resolved by a third, senior annotator who made the final adjudication. This process minimizes individual bias and errors. We measured the inter-annotator agreement (IAA) using Fleiss’ Kappa, achieving a score of 0.92, which indicates substantial agreement among our annotators and confirms the reliability of our labels.
N.2 Experimental Setup
@@
We evaluate our function-driven user simulator against a baseline simulator that directly generates responses using a single-pass LLM prompt. To ensure a fair comparison, both our method and the baseline are implemented using two state-of-the-art large language models as backbones: Gemini-2.0-Flash and GPT-4o.
@@
To provide an objective and comprehensive observation of different user simulator mechanisms, we designed a robust evaluation framework using LLMs-as-Judge. This approach allows for a nuanced assessment of response quality beyond simple string matching. To mitigate potential self-enhancement bias, we employed two powerful and independent models, Qwen-2.5-72B and Llama-3.1-70B, as evaluators.
@@
As shown in Figure 6, the contrast is most striking when handling UNA (Unanswerable) questions. Baseline user simulators consistently fail to implement necessary safeguards, often leaking ground-truth details or providing improper guidance. This leads to a high failure rate of over 34% for both model backbones. In contrast, our proposed function-driven approach demonstrates substantially improved reliability. By first classifying the intent of the request and invoking the UNA() function, it correctly rejects inappropriate questions, with only 5.9% of its responses falling into problematic categories. This represents a significant improvement in user simulator robustness.
Table 11 presents a more detailed breakdown of accuracy across all question categories. We observe that LLMs themselves already perform well on the first two categories (AMB and LOC), achieving over 90% accuracy even with the baseline approach. However, they struggle significantly with UNA (unanswerable) questions, where the baseline simulators fail in over 34% of cases for both backbone models. In contrast, our function-driven approach substantially mitigates this weakness, maintaining over 93% accuracy on UNA questions, which is a great improvement that confirms the observations from Figure 6. This demonstrates that while LLMs can naturally handle straightforward clarification tasks, they require explicit structural constraints to avoid inappropriately answering questions that should be refused. Our two-stage design enforces such constraints by first identifying the question type before generating a response, ensuring the simulator’s behavior remains predictable, controllable, and aligned with the goal of providing fair and realistic user feedback without leaking ground-truth information.
Table 11: Accuracy (%) of user simulators on the UserSim-Guard benchmark. Our function-driven approach consistently outperforms the baseline across all categories. Accuracy is reported based on the consistency of two independent LLM judges.
@@
Gemini-2.0-Flash Baseline 93.57 95.52 66.93
Ours (Function-Driven) 94.76 96.75 93.80
GPT-4o Baseline 90.72 96.30 63.45
Ours (Function-Driven) 98.81 98.88 94.08
@@
Appendix O Pathways to Effective Communication
@@
Appendix P Human Evaluation of Dataset Quality
@@
Appendix Q Prompts
Q.1 System Prompts
@@
Q.2 User Simulator Prompts

### v3

@@ Acknowledgement
Reynold Cheng, Nan Huo, Xiaohan Xu, Jinyang Li, Xiaolong Li, and Ge Qu were supported by the Research Grant Council of Hong Kong (RGC Project HKU 17202325), the University of Hong Kong (Project 2409100399), and the HKU Faculty Exchange Award 2024 (Faculty of Engineering). And we would like to express our sincere gratitude to Irina Saparina, Mohammadreza Pourreza, Mehdi Bouzouina, Hailong Li, Jiatong Shi, and Professor Shinji Watanabe for their fruitful discussions and valuable insights that helped improve this work.
Ethics Statement
This research complies with the ICLR Code of Ethics. We have carefully reviewed the guidelines and ensured that our work aligns with the stated ethical standards, including considerations of data privacy, fairness, and responsible use of released datasets and code. Justification: This work does not involve crowdsourcing or research with human subjects. All annotation and task creation were carried out by the authors themselves.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. First, for the evaluated systems, all experimental settings, including model parameters, temperature (=0=0), and budget configurations, are clearly documented in Appendix I, ensuring that our evaluation can be replicated under the same conditions. Each task is executed in a freshly re-initialized PostgreSQL 14 instance (Docker). The Docker image contains the database engine and benchmark code, and is restarted for every run to guarantee a clean and consistent state. This setup makes experiments deterministic, isolated across runs, and easy to reproduce by rebuilding the environment from scratch. Second, for the user simulator, we validate its robustness in Section 6, and detail in Section 3 how we designed and annotated the benchmark to guarantee reliability, uniqueness of clarifications, and consistency of responses from the user simulator. This safeguards reproducibility across different runs and systems. Third, for the benchmark suite, we will publicly release all components under a permissive license, including databases, tasks, hierarchical knowledge bases, documentation, interaction logs, and the source code for both evaluation settings and the user simulator. This full release ensures transparency and faithful replication of our experiments. We also provide the prompts used across the whole experiment, which can be found in Appendix R, from Figure 19 to Figure 26. Due to the dynamic nature of our interaction evaluation, we will also open-source our interaction trajectories upon publication for better reproducibility.
Experiment Configuration.
All experiments on Bird-Interact are conducted via API, except for the LLM-as-Judge evaluations of different user simulators, which are run on 4 NVIDIA A100 80G GPUs. The estimated cost for each model under Bird-Interact-Full is shown in Table 2, and the estimated cost for Bird-Interact-Lite is shown in Table 10.
@@ Appendix A Limitations
Appendix B The Use of LLM Statement
Large Language Models (LLMs) were used only for light post-editing of the paper (i.e., reducing syntax errors and performing minor grammar checks). LLMs were not involved in any part of the research discussions, analyses, or idea generation. All insights, contributions, and intellectual content are entirely the authors’ own.
Appendix C Annotation Group Details
@@
C.1 Annotator Entrance Test
@@ C.3 Qualification Test
Appendix D Benchmark Design Principles
@@
Appendix E Comparison with Related Benchmarks
E.1 Task Comparison
@@ E.2 Database Comparison
Appendix F Evaluation Metrics
F.1 Success Rate (SR)
@@
where $\mathbb{I}[\cdot]$ is the indicator function that equals 1 if the prediction is correct and 0 otherwise. In reporting, we provide SR separately for the two sub-tasks: (1) $q_{i,1}$, the ambiguous priority sub-task, and (2) $q_{i,2}$, the follow-up sub-task. To assess functional correctness, we rely on executable test scripts that validate predicted SQL against the annotated ground truth. Details of the test scripts are provided in Appendix G.
F.2 Normalized Reward
@@
Appendix G Test Scripts
@@ G.2 DM Queries
Appendix H Ambiguity and Follow-up Annotation Details
H.1 User Query Ambiguity Annotation
@@
For each category, we provide annotators with a structured taxonomy including type definitions, annotation conditions, and examples, ensuring systematic and consistent ambiguity injection, as outlined in Appendix H.4.
@@
H.2 Knowledge and Environmental Ambiguity Annotation
@@ H.3 Ambiguity Chain
H.4 User Query Ambiguity Taxonomy
@@
Intent-level ambiguity arises directly from user language, where the request is vague, underspecified, or missing critical details (e.g., “find elderly people” without defining the age threshold). If not resolved, intent-level ambiguity can severely degrade user experience and lead to erroneous SQL (Saparina & Lapata, 2024; Li et al., 2025f). We summarize six types of user query ambiguity in Table 6, according related works (Saparina & Lapata, 2024; Li et al., 2025f; Dong et al., 2025; Wang et al., 2020; Min et al., 2020; Bhaskar et al., 2023; Huo et al., 2024; Floratou et al., 2024; Huang et al., 2023; Ding et al., 2025; Xu et al., 2024b; a) and guide the annotators to inject them into unambiguous user queries: (1) Lexical Ambiguity from tokens with multiple meanings, (2) Syntactic Ambiguity from multiple valid grammatical structures, (3) Semantic Ambiguity from vague phrasing (e.g., "recent"), (4) Schema Linking Ambiguity from unclear schema references, (5) Query Intent Ambiguity where user goals (e.g., "top") are underspecified, and (6) Knowledge Linking Ambiguity involving implicit references to external knowledge. The performance of different types is shown in Figure 14.
@@
H.5 Follow-up Sub-Task Taxonomy
@@
Appendix I Experiment Details
I.1 Choice of PostgreSQL as the Evaluation Database System
@@ I.3 Experiment Setup
Appendix J Action Space and Selection Patterns in $a$-Interact
@@
J.1 Action Space in $a$-Interact
@@
J.2 Universal Cost Scheme for Custom Agents
@@
J.3 Action Selection Patterns and Their Impact (Full Set)
@@
Appendix K Performance on Different Ambiguity Types
@@
Appendix L Experiments on Bird-Interact-Lite
@@
Model Priority Question (Success Rate %) $\uparrow$ Follow Ups (Success Rate %) $\uparrow$ Reward* $\uparrow$ Avg. Cost $\downarrow$
@@
Gemini-2.0-Flash 8.21 44.76 21.00 4.10 21.90 10.33 17.80 \$ 0.03
DeepSeek-R1 6.67 47.62 21.00 3.59 28.57 12.33 18.40 \$ 0.09
GPT-4o 12.31 43.81 23.33 4.62 17.14 9.00 19.03 \$ 0.46
DeepSeek-V3 11.79 44.76 23.33 6.15 16.19 9.67 19.23 \$ 0.06
Qwen-3 7.18 49.52 22.00 5.64 29.52 14.00 19.60 \$ 0.03
O3-Mini 14.87 45.71 25.67 6.67 21.90 12.00 21.57 \$ 0.08
Claude-Sonnet-3.7 22.05 56.19 34.00 10.77 30.48 17.67 29.10 \$ 0.67
@@
Appendix M Error Analysis
@@
Appendix N User Simulator Design Details
@@
Appendix O Evaluating the Function-Driven User Simulator
@@
O.1 UserSim-Guard: A Benchmark for Simulator Robustness
To enable a systematic evaluation of simulator performance, we constructed UserSim-Guard, a manually curated dataset containing 2,100 challenging questions.
@@
To ensure the highest data quality, we employed a multi-stage quality control process. Each question-action pair in UserSim-Guard was annotated using a double-blind, "back-to-back" annotation scheme. Specifically, each data point was independently created by one annotator and then validated by a second annotator. Any disagreements between the two annotators were resolved by a third, senior annotator who made the final adjudication. This process minimizes individual bias and errors. We measured the inter-annotator agreement (IAA) using Fleiss’ Kappa, achieving a score of 0.93, which indicates substantial agreement among our annotators and confirms the reliability of our labels.
O.2 Experimental Setup
@@
We evaluate our function-driven user simulator against a baseline simulator that directly generates responses using a single-pass LLM prompt. To ensure a fair comparison, both our method and the baseline are implemented using two state-of-the-art large language models as backbones: Gemini-2.0-Flash, GPT-4o and Claude-Haiku-4.5.
@@
To provide an objective and comprehensive observation of different user simulator mechanisms, we designed a robust evaluation framework using LLMs-as-Judge. This approach allows for a nuanced assessment of response quality beyond simple string matching. We employed the Qwen3-235B-A22B-Instruct-2507 as evaluator.
@@
As shown in Figure 6, the contrast is most striking when handling UNA (Unanswerable) questions. Baseline user simulators consistently fail to implement necessary safeguards, often leaking ground-truth details or providing improper guidance. This leads to extremely high failure rates, reaching up to 67.4% depending on the backbone model. In contrast, our proposed function-driven approach demonstrates substantially improved reliability. By first classifying the intent of the request and invoking the UNA() function, it correctly rejects inappropriate questions, reducing the failure rate to as low as 2.7%. This represents a dramatic improvement in user simulator robustness.
Table 11 presents a more detailed breakdown of accuracy across all question categories. We observe that LLMs themselves already perform strongly on the ambiguity-related categories, achieving between 87.7% and 92.3% accuracy with the baseline approach. However, they struggle significantly with UNA (unanswerable) questions, where baseline accuracy drops as low as 32.6% for some backbones. In contrast, our function-driven approach substantially mitigates this weakness, consistently achieving between 87.3% and 97.3% accuracy on UNA questions. These improvements confirm the observations from Figure 6. This demonstrates that while LLMs can naturally handle straightforward clarification tasks, they require explicit structural constraints to avoid inappropriately answering questions that should be refused. Our two-stage design enforces such constraints by first identifying the question type before generating a response, ensuring the simulator’s behavior remains predictable, controllable, and aligned with the goal of providing fair and realistic user feedback without leaking ground-truth information. ***We consider a user simulator to be reliable if it achieves at least 90% accuracy on both labeled and unlabeled ambiguity categories, and at least 80% accuracy on unanswerable (UNA) questions.
Table 11: Accuracy (%) of user simulators on the UserSim-Guard benchmark using updated evaluation results. Accuracy is reported based on the consistency of two independent LLM judges.
@@
Gemini-2.0-Flash Baseline 89.6 89.3 40.3
Ours (Function-Driven) 94.9 93.6 87.3
GPT-4o Baseline 87.7 89.4 77.3
Ours (Function-Driven) 94.6 94.1 97.3
Claude-Haiku-4.5 Baseline 92.3 92.0 32.6
Ours (Function-Driven) 95.0 95.0 95.4
@@
Appendix P Pathways to Effective Communication
@@
Appendix Q Human Evaluation of Dataset Quality
@@
Appendix R Prompts
R.1 System Prompts
@@
R.2 User Simulator Prompts